OpenStudy (anonymous):

find the solution and equation , square root of x + 6 = 5

OpenStudy (anonymous):

\[\sqrt{x + 6}=5\] like this

OpenStudy (anonymous):

You would need to set it equal to 0, I think... Like \[\sqrt{x+6}-5=0\]

OpenStudy (anonymous):

\[\sqrt{A}=5 \rightarrow A=25\]

OpenStudy (anonymous):

That too

OpenStudy (anonymous):

So, then, building off Herman, take 25, and subtract 6.

OpenStudy (anonymous):

So your solution would be 19.

OpenStudy (anonymous):

i mean \[\sqrt{x+6}=5\]

OpenStudy (anonymous):

Yes. \[\sqrt{X-6}=5\rightarrow 5^2 = X-6\rightarrow X-6 = 25\] Solve for X

OpenStudy (anonymous):

Also, I goofed. The solution would in fact be 31.